OpenStudy (anonymous):

solve C = 5 over 9 (F - 32) for F

OpenStudy (mathstudent55):

Solve the equation for F \(C = \dfrac{5}{9}(F - 32) \)

OpenStudy (mathstudent55):

Solving for F means to isolate F. You need F by itself.

OpenStudy (mathstudent55):

The first thing to do is to get rid of the 5/9 fraction on the right side. To do that, multiply both sides by the reciprocal of 5/9 which is 9/5. Do that and show what you get.

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@kaos_gabz

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Part A: Solve C = 5 over 9 (F - 32) for F. (4 points) Part B: Determine the value of F when C = 8 °C. (2 points) Part C: Solve –np – 80 > 60 for n. Show your work. (4 points)

OpenStudy (anonymous):

i need all parts

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@marissalovescats

OpenStudy (anonymous):

@agent0smith

OpenStudy (mathstudent55):

Ok, I'm back.

OpenStudy (mathstudent55):

Part A: Solve C = 5 over 9 (F - 32) for F. (4 points) \(C = \dfrac{5}{9}(F - 32) \) The first step is to get rid of the 5/9, so we multiply both sides by 9/5: \(\color{red}{\dfrac{9}{5} \cdot} C = \color{red}{\dfrac{9}{5} \cdot}\dfrac{5}{9}(F - 32) \) \(\dfrac{9}{5} \cdot C = \dfrac{\color{red}{\cancel{9}}}{ \color{blue}{\cancel{5}} } \cdot\dfrac{\color{blue}{\cancel{5}}}{\color{red}{\cancel{9}}}(F - 32) \) \(\dfrac{9}{5} \cdot C = F - 32 \) Now we see that the only thing still with the F that we want to isolate is -32. 32 is being subtracted from F, so we add 32 to both sides. \(\dfrac{9}{5} \cdot C \color{red}{+ 32} = F - 32\color{red}{+ 32} \) \(\dfrac{9}{5} \cdot C + 32 = F \) \(F = \dfrac{9}{5} \cdot C + 32\)